Pre-Lab Assignment for Experiment 11: The frequency and the half wavelength d of a standing wave on a string are related by =tension F (N) 0.0 0.05 0.100.150.200.250.300.350.40 d? (m)
Answer #1

We have the expression relatinfg frequency and half wave length of the string related by,

f=\frac{1}{2d}\sqrt{\frac{F}{\mu }}

Where F =tension in the string in Newtons(N)

And,\mu=Linear mass density(kg/m)

Also from the equation,on squaring both sides,

f^{2}=\frac{F}{4d^{2}\mu }

Or ,ie F = (4uf?)*

So,From the table of the observation of the experiment,given in question,we find the different values of d^{2} by varying the tension in the string F(N) .

On plotting the graph of  d^{2}(m^{2}) vs F(N) :from the given values in the table,with d^{2}(m^{2}) on x axis and F(N) on y axis.

C D E F G H I J B 1 d2(m2) FIN) 2 0.325 5.47 3 0.179 3.04 0.114 1.96 0.08 1.34 d2(m2) v/s F(N) y = 16.781x +0.0243 5.47 F(N)

Here the equation of the line of fit found to be,as shown,y=16.781x+0.0243

It is the equation of the line,y=mx+c

So,the slope of the line fit,m=16.781N/m^{2}

So,here we have the slope of the line from the equation that,m=\frac{F}{d^{2}}=4\mu f^{2}= 16.781N/m^{2}

So,we have 4\mu f^{2}= 16.781N/m^{2}

Also we have the slope,m=\frac{F}{d^{2}}

For finding the uncertainity in slope,relative uncertainity in slope,\frac{\Delta m}{m}=\frac{\Delta F}{F}+\frac{\Delta (d^{2})}{d^{2}}

Here,the mean value of d^{2}=\frac{0.325+0.179+0.114+0.08}{4}=0.1745m^{2}

And the mean value of \Delta d^{2}=\frac{0.017+0.013+0.01+0.009}{4}=0.01225m^{2}

mean value of \Delta F=\frac{0.34+0.10+0.10+0.03}{4}=0.1425N

Mean value of F=\frac{5.47+3.04+1.96+1.34}{4}=2.9525N

So,ie,\frac{\Delta m}{m}=\frac{0.1425}{2.9525}+\frac{0.01225}{0.1745}=0.1185

Also we have from the graph,slope,m=16.781N/m^{2}

So,eqn becomes,\Delta m=0.1185*m=0.1185*16.781=1.9885N/m^{2}

uncertainity of slope,\Delta m=1.9885N/m^{2}

So,the slope of the graph with uncertainity,m\pm \Delta m=16.781\pm 1.9885N/m^{2}

We have given for the value of linear mass density,\mu =0.315g/m^{2}=3.15*10^{-4}kg/m^{2}

So,for this value of \mu ,the value of frequency can be find from slope,m=4\mu f^{2}

So,4\mu f^{2}=m=16.781N/m^{2}

Then,frequency,f=\sqrt{\frac{m}{4\mu }}=\sqrt{\frac{16.781}{4*3.15*10^{-4}}}=115.404Hz

So,frequency,f=115.404Hz

Here to find uncertainity in frequency,\frac{\Delta m}{m}=2*\frac{\Delta f}{f}

So,\Delta f=\frac{\Delta m*f}{2m}=\frac{1.9885*115.404}{2*16.781}=6.837Hz

And the frequency,f\pm \Delta f=115.40\pm 6.837Hz

Or the value of frequency,f=115.404Hz

Please upvote the answer...